OpenStudy (anonymous):

simplify the expression: (5s^-2t^4)^-3

OpenStudy (anonymous):

\[ \frac{s^6}{125 t^{12}} \]

OpenStudy (anonymous):

(1/125)*s^6*t^12

jhonyy9 (jhonyy9):

1/((5t4)/s2)3 = 1/125t12/s6 = s6/125t12